Publications of the Syracuse University Computational Physics Group

1. C.F. Baillie and P.D. Coddington, Comparison of Cluster Algorithms for 2-D Potts Models, Phys. Rev. B 43, 10617 (1991). SCCS-12.

2. J. Apostolakis, C. Baillie and G.C. Fox, Investigation of the 2d O(3) Model Using the Overrelaxation Algorithm, Phys. Rev. D 43, 2687 (1991). SCCS-65.

3. J. Apostolakis, C. Baillie and G.C. Fox, The 2d O(3) Model by Overrelaxation, Proc. of the International Conference on Lattice Field Theory, Tallahassee (October 1990), Nucl. Phys. B (Proc. Suppl.) 20, 678 (1991). SCCS-101.

4. P.D. Coddington and C.F. Baillie, Parallel Cluster Algorithms, Proc. of the International Conference on Lattice Field Theory, Tallahassee (October 1990), Nucl. Phys. B (Proc. Suppl.) 20, 76 (1991). SCCS-102.

5. J. Apostolakis, P. Coddington and E. Marinari, A Multi-Grid Cluster Labeling Scheme, Europhys. Lett. 17, 189 (1992). SCCS-103.

6. A.A. Middleton, Thermal Rounding of the Charge Density Wave Depinning Transition, Phys. Rev. Lett. 45, 9465 (1992). SCCS-108.

7. P.D. Coddington and C.F. Baillie, Empirical relations between static and dynamic exponents for Ising model cluster algorithms, Phys. Rev. Lett. 68, 962 (1992). SCCS-141.

8. D.J. Edelsohn, Hierarchical Tree-Structures as Adaptive Meshes, to be published in Int. J. Mod. Phys. C. SCCS-193.

9. A.A. Middleton, O. Biham, P.B. Littlewood and P. Sibani, Complete Mode Locking in Models of Charge Density Waves, Phys. Rev. Lett. 68, 1586 (1992). SCCS-217.

10. P.D. Coddington and C.F. Baillie, Dynamical Exponents for Potts Model Cluster Algorithms, Proc. of the International Conference on Lattice Field Theory, Tsukuba, Japan (November 1991), Nucl. Phys. B (Proc. Suppl.) 26, 632 (1992). SCCS-220.

11. E. Marinari and G. Parisi, Simulated Tempering: A New Monte Carlo Scheme, Europhys. Lett. 19, 451 (1992). SCCS-241.

12. G. Iori, E. Marinari, G. Parisi and M.V. Struglia, Statistical Mechanics of Heteropolymer Folding, Physica A185, 98 (1992).

13. M. Guagnelli, M.P. Lombardo, E. Marinari, G. Parisi and G. Salina, The Quenched Mass Spectrum in Lattice QCD, Proc. of the International Conference on Lattice Field Theory, Tsukuba, Japan (November 1991), Nucl. Phys. B (Proc. Suppl.) 26, 278 (1992).

14. M. Guagnelli, M.P. Lombardo, E. Marinari, G. Parisi and G. Salina, The Quenched Mass Spectrum in Lattice QCD on a 1 Gigaflops Computer, Nucl. Phys. B378, 616 (1992).

15. E. Marinari, A. Middleton, Statistical Mechanics on Massively Parallel Computers: The Research at NPAC, Parallel Computing News 5, 21 (1992). SCCS-452.

16. T. Haupt, Visualization of High Energy Physics Data Using AVS. SCCS-243.

17. A. Pertsov, M. Vinson, and S. Müller, Three-dimensional reconstruction of organizing centers in excitable chemical media, Physica D 63, 233 (1993). SCCS-247.

18. E. Marinari et al., The APE-100 Computer: The Architecture, International Journal of High Speed Computing, to appear. SCCS-268.

19. M. Bowick and E. Marinari, Quantum Gravity, Random Geometry and Critical Phenomena, Gen. Rel. Grav. 24, 1209 (1992), (Fourth Award in the Gravity Research Foundation Essay Competition 1992). SCCS-269, SU-4241-505.

20. J. Apostolakis, P. Coddington and E. Marinari, New SIMD Algorithms for Cluster Labeling on Parallel Computers, Int. J. Mod. Phys. C 4, 749 (1993). SCCS-279.

21. M. Guagnelli, E. Marinari and G. Parisi, Mean Field Solutions of the Random Field Ising Model, J. Phys. A (Math. Gen.), to appear. SCCS-329.

22. P. Pliszka and E. Marinari, On Heteropolymer Shape Dynamics, Europhysics Lett. 22, 167 (1993). SCCS-330.

23. G. Iori, E. Marinari and G. Parisi, Non-Exponential Relaxation Time Scales in Disordered Systems: an Application to Protein Dynamics, submitted to J. Phys. A. (Math. Gen.) SCCS-331, ROM2F-92-35.

24. M. Dong, M.C. Marchetti, A.A. Middleton and V. Vinokur, Elastic String in a Random Potential, Phys. Rev. Lett. 70, 662 (1993). SCCS-346.

25. D.S. Fisher and A.A. Middleton, Critical Behavior of Charge Density Waves Below Threshold: Numerical and Scaling Analysis, Phys. Rev. B. 47, 3530 (1993). SCCS-347.

26. O. Biham, A.A. Middleton, and D. Levine, Self-Organization and a Dynamical Transition in Traffic Flow Models, Phys. Rev. A 46, R6124 (1992). SCCS-348.

27. M. Bowick, P. Coddington, L. Han, G. Harris and E. Marinari, The Phase Diagram of Fluid Random Surfaces with Extrinsic Curvature, Nucl. Phys. B 394, 791 (1993). SCCS-357, SU-HEP-4241-517.

28. B. Brügmann, Non-uniform Measure in 4d Simplicial Quantum Gravity, Phys. Rev. D 47, 3330 (1992). SU-GP-92/9-1.

29. B. Brügmann and E. Marinari, 4D Simplicial Quantum Gravity with a Nontrivial Measure, Phys. Rev. Lett. 70, 1908 (1993). SCCS-361, SU-GP-92/9-2.

30. M.L. Paciello, E. Marinari, G. Parisi and B. Taglienti, The String Tension in Gauge Theories: a Suggestion for a New Measurement Method, Phys. Lett. 298B, 400 (1993).

31. O. Biham and M. Kuale, Unstable Periodic Orbits in the Stadium Billiard, Phys. Rev. A 46, 6334 (1992).

32. E. Marinari, M. Guagnelli, and G. Parisi, Maximal Mean Field Solutions in the Random Field Ising Model: the Pattern of the Symmetry Breaking, submitted to J. Phys. A. SCCS-372, ROM2F-92-46, cond-mat/9303042.

33. M. Bowick, P. Coddington, L. Han, G. Harris and E. Marinari, Strings with Extrinsic Curvature: An Analysis of the Crossover Regime, Proc. of the International Conference on Lattice Field Theory, Amsterdam (October 1992), Nucl. Phys. B (Proc. Suppl.) 30, 795 (1993). SCCS-401, SU-HEP-4241-527.

34. G. Iori, E. Marinari and G. Parisi, Random self-interacting chains: a mechanism for protein folding, J. Phys. A (Math. Gen.) 24, 5349 (1991). SCCS-382.

35. E. Marinari and G. Parisi, On Polymers with Long Range Repulsive Forces, Europhys. Lett. 15 (1991) 721.

36. E. Marinari, A Review Talk about Computers and Theoretical Physics, Proc. of the International Conference on Lattice Field Theory, Amsterdam (October 1992), Nucl. Phys. B (Proc. Suppl.) 30, 122 (1993). SCCS-383.

37. M. Bowick, B. Brügmann, P. Coddington, L. Han, G. Harris and E. Marinari, Random Surfaces and Quantum Gravity: Looking for the Emergence of Continuum Theories from Triangulated Manifolds, Proceedings of the Conference “String Theory, Quantum Gravity and the Unification of the Fundamental Interactions”, eds. M. Bianchi, F. Fucito, E. Marinari and A. Sagnotti, (World Scientific, Singapore, 1993). SCCS-387.

38. M. Bowick, P. Coddington, L. Han, G. Harris and E. Marinari, The Phase Structure of Strings with Extrinsic Curvature, Proc. of the Workshop “String Quantum Gravity and Physics at the Planck Energy Scale”, Erice, Italy (June 1992). SCCS-391, SU-HEP-4241-528.

39. G. Fox, J. Apostolakis, M. Bowick, B. Brügmann, P. Coddington, D. Edelsohn, L. Han, G. Harris, E. Marinari, K. Mills and S. Ranka, Computer Science Software Tools and Languages to Support Computational Physics, report to the IBM Corporation. SCCS-393.

40. M. Vinson, A. Pertsov and J. Jalife, Anchoring of vortex filaments in 3D excitable media, submitted to Physica D.

41. P. Ensminger and M. Vinson, Deterministic nonlinear dynamics in Phycomyces sporangiophores, submitted to Journal of Theoretical Biology.

42. M. Bowick, P. Coddingtion, G. Fox, L. Han, G. Harris, and E. Marinari, Optimization of a Dynamic Random Surface Program for RISC Processors, presented at the SIAM 1993 Meeting, Philadelphia, PA (July 1993). SCCS-481.

43. V. Dotsenko, P. Windey, G. Harris, E. Marinari, E. Martinec and M. Picco, Critical and Topological Properties of Cluster Boundaries in the 3d Ising Model, Phys. Rev. Lett. 71, 811 (1993). SCCS-621, SU-HEP-4240-531

44. E. Marinari and G. Parisi, On Toy Aging, J. Phys. A (Math. Gen.), to appear.

45. Zheng-Yao Su, Wen Chen and Geoffrey C. Fox, The Implementations of the Hierarchical Cluster Labeling Algorithm on Parallel Computation Environments, to appear in Proc. 2nd Int. Conf. on Computational Physics, Beijing (1993).

46. O. Biham, L.W. Chen and G. Vidali, Models for monolayers adsorbed on a square substrate, Surface Science 287, 815 (1993).

47. W. Li, G. Vidali and O. Biham, Scaling of island growth in Pb overlayers on Cu(001), Phys. Rev. B, to appear.

48. G. Barkema, O. Biham, M. Breeman, D.O. Boerma and G. Vidali, Characterization of submonolayer growth of Cu Islands on Cu(001), submitted to Phys. Rev. Lett.

49. G. Vidali, O. Biham, H. Zeng, J.S. Lin and W. Li, Scaling in the growth of Pb islands on Cu(001), Proc. of the 4th Int. Conf. on the Structure of Surfaces, to appear (1993).

50. P.D. Coddington, An Analysis of Distributed Computing Software and Hardware for Applications in Computational Physics, to be published in the Proceedings of the Second International Symposium on High Performance Distributed Computing, Spokane, WA (July 1993). SCCS-421.

51. G.C. Fox and P.D. Coddington, An Overview of High Performance Computing for the Physical Sciences, to be published in the Proceedings of the Mardi Gras Conference on High Performance Computing and Its Applications in the Physical Sciences, Baton Rouge, LA (February 1993), ed. D.A. Browne et al., (World Scientific, Singapore, 1994). SCCS-488.

52. K. Anagnostopoulos, M. Bowick, P. Coddington, M. Falcioni, L. Han, G. Harris and E. Marinari, Fluid Random Surfaces with Extrinsic Curvature: II, Phys. Lett. B 317, 102 (1993). SCCS-511, SU-HEP-93-4241-540.

53. J. Apostolakis, Asymptotic scaling in the two-dimensional nonlinear Sigma model: A Monte Carlo study on parallel computers, Ph.D. thesis. SCCS-514.

54. Z.-Y. Su, The Hierarchical Algorithms – Theory and Applications, Ph.D. thesis. SCCS-524.

55. P.D. Coddington, Analysis of Random Number Generators Using Monte Carlo Simulation, Int. J. Mod. Phys. C 5, 547 (1994). SCCS-526.

56. P.D. Coddington and L. Han, Generalized Cluster Algorithms for Frustrated Spin Models, Phys. Rev. B. 50, 3058 (1994). SCCS-527.

57. P.D. Coddington and L. Han, A Generalized Cluster Algorithm for the 3-d Fully Frustrated Ising Model, to be submitted to Phys. Rev. B. SCCS-528.

58. E. Marinari, G. Parisi and F. Ritort, On the 3d Spin Glass, J. Phys. A (Math. Gen.), to appear. SCCS-532.

59. M. Bowick, M. Falcioni, G. Harris and E. Marinari, Two Ising Models Coupled to 2–Dimensional Gravity, Phys. Lett. B, to appear. SCCS-540, SU-HEP-93-4241-556.

60. G. Harris, Percolation on Strings and the Cover-up of the c = 1 Disaster, SU-HEP-4241-555.

61. M. Bowick, M. Falcioni, G. Harris and E. Marinari, Critical Slowing Down of Cluster Algorithms for Ising Models Coupled to 2-d Gravity, Phys. Lett. B 322, 316 (1994). SCCS-568, SU-HEP-4241-560.

62. S.-J. Bae, S.H. Ko and P.D. Coddington, Parallel Wolff Cluster Algorithms, submitted to Int. J. Mod. Phys. C. SCCS-619.

63. M. Bowick, G. Harris, M. Falcioni and E. Marinari, Multiple Ising Models Coupled to 2-d Gravity: A CSD Analysis, Proc. of the International Conference on Lattice Field Theory, Dallas (Oct. 1993).

64. G.C. Fox and P. D. Coddington, Parallel Computers and Complex Systems, in T. Bossomaier and D. G. Green, editors, Complex Systems, (Cambridge University Press, Cambridge, in press).